OpenStudy (anonymous):

A painter leans a 34-ft ladder against a building. The base of the ladder is 16 ft from the building. To the nearest foot, how high on the building does the ladder reach?

OpenStudy (anonymous):

c^2=a^2+b^2 34^2=16^2+b^2 900=b^2 b=30ft

OpenStudy (anonymous):

Thank you! @coolaserman